In Millikan's experiment, the charge on each drop of oil was measured in coulombs. Imagine the same experiment, but with charges measured in a fictitious unit called a zeet (Z ).

Drop Charge (Z )
A 7.80×10−15
B 1.82×10−14
C 2.08×10−14
D 2.34×10−14
E 3.38×10−14

What is the charge on an electron in zeets?

Express your answer in zeets using two significant figures.
Answer #1

Electron is indivisible and carries specific charge. We first look at the numbers representing the charges in zeets and try to find the largest common factor.

2.08 - 1.82 = 0.26

2.34 - 2.08 = 0.26

0.26 is found to be lowest difference between two figures. Test this as follow

78/0.26 = 300

1.82/0.26 = 7

2.08/0.26 = 8

2.34/0.26 = 2.08

3.38/0.26 = 13

In all the above cases, quotients are integers, so it can be said that these quotients are the actual number of excess electrons on the drops, in which case the charge on an electron is 0.26 x 10-14 zeet.